OpenStudy (anonymous):

the sqaure root of 2 is a rational number true or false

OpenStudy (lgbasallote):

can you express sqrt 2 as a quotient of two INTEGERS?

OpenStudy (ash2326):

@heartbreak do you know what's a rational number?

OpenStudy (anonymous):

The square root of 2, often known as root 2 or radical 2 and written as , is the positive algebraic number that, when multiplied by itself, gives the number 2. It is more precisely called the principal square root of 2, to distinguish it from the negative number with the same property.

OpenStudy (lgbasallote):

..lol...

OpenStudy (anonymous):

Geometrically the square root of 2 is the length of a diagonal across a square with sides of one unit of length; this follows from the Pythagorean theorem. It was probably the first number known to be irrational. Its numerical value truncated to 65 decimal places is: 1.41421356237309504880168872420969807856967187537694807317667973799... (sequence A002193 in OEIS) ucan see here http://en.wikipedia.org/wiki/Square_root_of_2

OpenStudy (ash2326):

@best.shakir I don't see, how this is answering the question here!

OpenStudy (anonymous):

ok so it's true @best.shakir

OpenStudy (anonymous):

i ment false

OpenStudy (ash2326):

Yeah @heartbreak do you know the reason?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

see down of the first link

OpenStudy (anonymous):

@ash2326 did u see

OpenStudy (ash2326):

@heartbreak a rational number is of the form \[\frac ab\] where a and b are positive or negative integers Some example of a rational no.= 2, 3, -1, 4/6, 7/8 etc. do you get this?

OpenStudy (anonymous):

yes i understand thank you

OpenStudy (ash2326):

now, if we try to find the root of 2, \[\sqrt 2=??\] it can't be expressed as a rational no.

OpenStudy (anonymous):

yes beacuse i did that and i got 1.41

OpenStudy (ash2326):

it's not just 1.41 but it'll have infinite digits after 1 too. that's why it can't be expressed as a/b had it been just 1.41=141/100 =rational

OpenStudy (anonymous):

o ok

OpenStudy (ash2326):

do you get my point?

OpenStudy (anonymous):

yes